A Complete Bibliography of Publications in Internet Histories

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 October 2025 Version 1.05

Title word cross-reference

$27.95 [169]. $30 [214]. $89 [214].

/dev/null [15].

0 [230].

19 [216, 208]. 1916 [48]. 1950s [104]. 1980s [104]. 1990 [91]. 1990s [222, 72, 115, 53, 236, 248]. 1990s-era [236].

2008 [149]. 2017 [60]. 2018 [89]. 2023 [245]. 220pp [232]. 25 [62]. 2channel [146].

3.0 [71].

4 [70, 232].

5 [42]. 5-Star [42]. 50/50 [227].

6 [179].

7 [231, 225, 224].

9 [189]. 90s [51]. 978 [231, 189, 225, 232, 224, 230]. 978-0-19-887597-0 [230]. 978-0-262-04751-7 [225]. 978-1-4798-1494-7 [231]. 978-1-4798-1831-0 [214]. 978-1-5036-3288-2 [224]. 978-1-5036-3704-7 [224]. 978-1-5095-4108-9 [189]. 978-1-5095-5052-4 [232].

Aaron [196]. abandoned [94]. Abidin [189]. abundance [102]. academia [194]. academic [98, 180]. access [175]. Accessing [216]. accordingly [16]. action [110]. activism [149, 201]. acts [201]. adoption [222, 186]. African [14]. Against [133]. Age [111, 222, 149, 28, 71, 244, 92, 102, 135, 112, 250, 225]. agenda [19]. AIDS [227]. Alan [49]. Aleksandra [129]. Aleph [171]. Aleph-bet [171]. Alexandre [67]. algorithmic [24, 100]. alt [145]. alt-right [145]. alternative [202, 55]. am [138]. ambiguity [40]. Amelia [232]. America [32, 52]. American [126]. Amsterdam [16]. Analysing [132]. analysis [107, 191, 218, 208]. analytics [128]. anarchist [240]. Anarchy [199]. Andrew [61]. Anglophone [56]. anime [177]. Ann [189]. années [91]. anonymity [164]. antecedents [119]. anxieties [166]. App [232]. approach [84, 211, 182]. approaches [134]. archaeological [182]. Archaeology [16, 76]. archival [134, 205]. archive [200, 48, 221, 46, 243, 47, 160, 17]. archived [181, 92]. Archives [179, 187, 39, 204, 244, 167, 83, 249]. archiving [84, 181, 180, 160]. Archivists [250]. Area [208]. Ariadna [232]. Armed [46]. ARPANET [44, 68, 65, 66, 69, 67, 64, 151]. arrangements [105]. art [81, 209, 208]. artificial [71]. artwork [216]. Arwid [50]. Asian [141, 151]. assets [108]. asymmetrical [11]. Australian [212, 248]. author [188]. authoritarianism [133]. authority [178]. authority-building [178]. Automatic [168]. autopsy [164]. Averting [250]. Avery [214]. award [113, 176, 226]. awful [191]. awry [164].

Back [67, 164]. Balbi [230]. band [77]. bans [165, 159]. Baran [27]. Barlow [172]. based [134, 119, 13]. Baulch [232]. Bay [208]. BBS [149]. be [16, 160]. became [65]. Becoming [235, 43]. before [105]. behavior [21]. Behind [57]. behold [33].

Beijing [149]. being [24, 71]. Benjamin [34]. Bernhard [100]. bet [171]. Between [184, 144, 201]. beyond [203]. Beyssac [57]. bias [131]. big [244]. biopolitics [159]. birth [200, 55]. Bit [61]. Blacksburg [162]. blasphemy [3]. Bloggernacle [238]. bloggers [238]. Blogging [70]. Bodies [118]. Book [21, 49, 231, 60, 91, 189, 28, 169, 89, 70, 225, 88, 123, 90, 232, 174, 224, 35, 214, 22, 34, 230]. Bory [123]. both [54]. boundary [66]. Brazil [116]. breakdown [240]. Breaking [18, 198]. breakup [194]. Brilliant [248]. Broad [77]. brought [248]. Brügger [90]. Building [37, 55, 238, 178]. Built [250]. Buzz [187].

C. [21]. Cambridge [169, 89, 174]. Can [5, 10, 160]. Capitalism [50, 111]. Capitol [179]. CAPTCHA [127]. Capturing [48]. Career [226, 113, 176]. Cascading [186]. case [193, 23, 40, 13]. cave [7]. Central [32, 52]. CERN [200]. cernes [204]. Ceruzzi [89]. challenges [108, 229, 13, 234]. changing [95, 117, 158, 185]. Chant [207]. character [13, 171]. character-based [13]. Chase [231]. Chinese [225, 149, 147, 150, 152]. Choking [148]. Chon [151]. Christian [56]. circle [150]. city [16, 182]. civil [114]. class [119]. class-based [119]. Classification [168]. close [249]. closed [95]. Club [138]. cluster [218]. co [41, 188]. co-author [188]. co-production [41]. code [85, 171]. coll [91]. collaboratory [31]. collection [200]. comes [135]. Commemorate [231]. commemoration [48]. comment [185]. commentary [128]. commercial [117]. commercialized [115]. common [172, 4]. Communication [232, 117, 120, 180, 135, 30]. Communications [88, 107]. communities [41, 237, 185, 184]. community [239, 145, 204]. comparison [75]. competing [201]. competition [104]. components [217]. Computational [80, 79, 78, 83]. Computer [225, 138, 194, 136, 52, 95]. computers [120]. computing [95, 7, 142]. Concealing [220]. conference [245]. confidential [21]. connected [33]. connecting [52]. consecration [172]. construction [68, 47, 200, 91]. constructivist [211]. contemporary [192]. Content [126, 227, 125]. Continuity [204]. Controverses [91]. controversy [99]. Conversation [44, 68, 81, 69, 151]. Conversations [27]. conversion [26]. Cookies [99, 105]. copyright [212, 126]. Corporation [139]. Correction [140]. Couchsurfing [158]. counterculture [58]. COVID [216, 208]. COVID-19 [216, 208]. cowboy [202]. craigslist [130, 75]. crawling [205]. creating [209]. creation [178]. Creative [152]. criminologist [38]. cross [84]. cross-platform [84]. Crypto [199, 29]. crypto-discourse [29]. Cryptoeconomics [199]. Crystal [189]. CSS [186]. cultural [74, 160, 5]. culture [196, 54]. Cultures [101, 81, 146, 58]. current [195]. curve [122]. cyberactivism [116]. cybercrime [38]. cyberlibertarian [172]. cyberlibertarianism [196]. cyberspace [56]. cynical [146]. Cypherpunk [173, 133].

Daemons [88]. dahs [171]. Dallto [210]. Dame [214]. Dame-Griff [214]. Danescu [110]. Danish [132, 218]. dans [91]. Dark [250, 251, 87]. data [16, 131, 244, 221]. Dating [183, 182]. days [44]. Dead [155, 154, 166]. Dead-and-dying [155, 154]. Death [231, 156, 159, 161, 167]. deaths [158]. debate [37]. debates [185]. decade [51]. decline [154]. Defence [46]. defending [55]. define [29]. defining [2]. democracy [42]. democratic [202]. Demonstrating [186, 18]. Departed [231]. describing [162]. design [68, 8, 186]. designed [131]. destruction [167]. dev [15]. Development [137, 11, 200, 73, 157, 46, 30]. developmental [144]. devices [66]. diachronic [218]. Diagnosing [156]. dialectics [195]. Digital [101, 231, 238, 146, 88, 248, 230, 220, 16, 222, 213, 246, 108, 95, 19, 147, 182, 237, 201, 202, 154, 25, 92, 55, 217, 12, 58, 197, 112, 152, 208, 250, 210, 21]. d’Internet [91]. direct [42]. directed [33]. discontinued [94]. discontinuity [204]. discourse [95, 29, 42]. discourses [177]. Discursive [149]. discussion [79, 109]. disembodiment [118]. Disqus [185]. dissolution [117]. distant [249]. dits [171]. dits-and-dahs [171]. divides [12]. Do [177]. doc [79, 109]. documentaries [17]. documenting [179]. does [156, 234]. Doing [17, 243, 92, 157]. Don’t [164, 118]. Driscoll [60]. during [149, 208]. Dutch [76]. Dwelling [244]. dying [155, 154]. dynamic [16, 19]. dynamics [178].

Early [206, 13, 44, 113, 176, 246, 118, 134, 144, 180, 119, 54, 56, 226]. earth [228]. East [11]. Easter [48]. ebook [224]. EC [104]. ecology [158]. economic [105]. economy [175]. ecosystems [184]. edited [169, 90]. Editorial [103, 45, 215]. efficacy [114]. Electronic [162, 143]. Elena [110]. Elephants [80]. Élie [69]. emergence [106]. Emigration [192]. Emma [232]. enablers [234]. enclaves [202]. encrypted [85]. Encryption [220, 29]. end [156, 85, 157, 122]. end-to-end [85]. enduring [202]. engineer [38, 138]. engines [24]. environmental [228]. epidemic [227]. epistemic [128]. epistemological [229]. era [236, 142, 234]. Ervin [61]. essay [119]. essence [223]. ethics [56]. Études [91]. EU [103, 96]. eulogies [167]. Euronet [96, 76]. Europe [108]. European [107, 106, 109]. evaluating [134]. Everything [160]. Evolution [86, 241, 31, 132, 73, 152]. examining [119]. exclusionary [177]. exes [161]. exhibition [208]. exist [177]. experimental [248]. Experimenting [83]. extremist [146].

Fabrique [91]. Face [231, 95]. Facebook [73]. facto} [85]. facts [47]. failure [154]. Fake [224]. fall [72, 163]. fan [237]. Fandom [101]. fandoms [159]. fans [53, 177]. features [198]. feeling [9]. Felsenstein [138]. female [53, 177]. feminist [244]. Fenwick [88]. Fernández [232]. fervent [145]. Fi [174]. fiction [237]. FidoNet [240]. field [110, 120, 135]. Fight [111]. Film [33]. filtering [126]. financialisation [10]. first [194, 62]. firsttuesday} [204]. Fixing [228]. flame [39]. flames [227]. Follow [221]. Football [101]. footer [209]. Force [235]. forced [26]. Forces [46]. Forensic [134]. Forgotten [161, 37, 142, 94]. found [84]. foundation [115]. four [158]. fragility [35]. frameworks [181]. Framing [213]. francaise [91]. France [55, 70]. fraught [119]. free [196, 138]. Freedom [50, 220, 196, 29]. Freenet [87]. French [121, 187, 69]. Friendster [156]. front [157]. front-end [157]. Frontier [111]. Future [251, 139, 111, 40, 42].

G [49]. Gabriele [230]. gallery [209]. Games [61]. gaming [239]. Gap [235]. Garden [186]. Geert [58]. gender [245, 241, 233, 242, 166, 235]. gendered [119]. Gendering [236]. genealogical [38]. Genealogies [125]. Genealogy [200]. genesis [192]. GeoCities [167]. Gérard [69]. Global [225, 232, 114, 97]. globalizing [148]. godfather [151]. gone [164, 161]. good [234]. Goodbye [121]. Google [137]. governance [212, 115, 114, 201, 199]. GPS [89]. Greater [208]. greatest [65]. Greek [222]. Griff [214]. Gross [49]. Guangdong [208]. Guy [32].

hacking [195]. Hagiography [3]. Han [13, 22]. handbook [90]. Hangul [144]. hardcover [224, 214]. Hargittai [21]. Harmon [49]. harvest [76]. Hauben [188]. Having [54]. Header [209]. Header/footer [209]. healthy [185]. heart [240]. Hebrew [171]. Hendry [189]. her [110]. Herding [80]. Herzog [33]. hidden [53]. historians [39]. historical [222, 47, 191, 102, 43, 100]. histories [2, 79, 78, 176, 226, 6, 10, 233, 141, 202, 3, 94, 243, 234, 14, 18, 62, 8, 113, 1]. History [225, 219, 102, 224, 214, 230, 245, 133, 246, 19, 190, 29, 59, 134, 53, 82, 203, 235, 198, 7, 127, 96, 97, 144, 37, 242, 25, 188, 92, 199, 116, 170, 12, 17, 67, 223, 5, 112, 34, 178, 90]. HIV [227]. HIV/AIDS [227]. Homebrew [138]. HOMOCORE [240]. homogenization [170]. Hong [208]. hope [175]. hopes [72]. Hot [65]. Hourtin [121]. HTML [98]. Hub [247]. Human [111, 24, 136, 71]. Humanities [49]. Hungarian [223]. hybrid [31].

I2P [87]. Ian [90]. idea [104]. ideal [236]. ideas [185]. identification [125]. Identities [101, 213]. identity [241, 105, 217]. ideological [24]. Ideologies [123]. ideology [172, 173, 228, 230]. If [139, 5]. Ilbe} [145]. imaginaire [203]. Imaginary [123, 148]. Imagining [180, 157]. Immortality [231]. impact [188, 177]. important [23]. IMPs [66]. Ina [91]. Incomplete [169]. incunables [76]. indentured [166]. India [148, 197]. inequality [175]. influence [181]. influences [173]. Infonie [26]. Information [225, 196, 110]. infrastructural [40, 205, 154]. infrastructure [193, 132, 181, 73]. infrastructures [66]. inherit [97]. Innovation [122]. Instagram [43]. instrumentalised [136]. intangible [108]. intellectual [133, 199]. intelligence [71]. inter [96, 66]. inter-structuralism [66]. interfaces [24]. Interfacing [58]. international [115, 207]. Internet [49, 60, 123, 226, 236, 2, 245, 11, 193, 23, 121, 47, 130, 79, 78, 176, 196, 40, 10, 103, 86, 132, 19, 233, 106, 141, 190, 115, 114, 29, 77, 134, 53, 82, 7, 120, 96, 97, 26, 144, 180, 37, 119, 13, 192, 88, 42, 160, 151, 148, 55, 135, 17, 3, 51, 67, 109, 32, 52, 143, 94, 9, 197, 5, 243, 234, 14, 18, 150, 152, 30, 224, 214, 34, 8, 1, 113, 188]. Interview [129, 207, 188, 138, 57, 58, 32, 210, 100, 110]. Intimacy [183]. Introduction [1, 78, 233, 125, 141]. Invented [139]. invention [181]. inversion [205]. Invisible [131]. Iperbole} [72]. IPO [137]. IPv6 [40]. ISBN [231, 189, 225, 232, 224, 214, 230, 70]. ISBN-978-1-4331-4710-4 [70]. issues [108]. Italian [72]. iWiW [223].

J [174, 224]. Jackie [169]. January [179]. Japan [146]. JARING [143]. job [234]. John [172, 28]. Johns [232]. Joseph [49, 169]. journalistic [172]. Julien [60]. justice [147, 243]. justice-seeking [147].

Kaminska [129]. Katrin [189]. Kevin [60]. key [190, 217]. Kilnam [151]. KISS [248]. Kleinrock [44]. Know [170]. knowledge [41, 178]. Koerner [169]. Kong [208]. Korea [145]. Korean [144].

labor [166]. labour [180]. Lang [70]. language [4, 56]. Lann [69]. large [83]. large-scale [83]. Larry [68, 54]. law [126, 104]. Laws [183]. Laying [115]. LCHC [31]. led [68]. Lee [138]. legacies [96]. Legacy [94, 98, 222, 65, 99]. Legitimacy [87, 190]. lens [212]. Leonard [44]. lessons [237, 37]. Let [191]. letter [53]. level [204]. LGBTQ [243]. Liberties [220]. Librarians [250]. library [181]. Life [231, 161, 71]. lifetime [166]. lineage [74]. list [39]. Lithuania [193]. Little [127]. local [97]. Long [161]. Long-Gone [161]. looking [39]. loss [216]. lost [96]. Love [183]. Lovink [58]. Lund [50].

M5S [42]. MA [169, 89]. Macao [208]. made [77]. magazine [118]. magazines [236]. magic [150]. Mailing [39]. mailinglist [81]. Mailland [60]. mainstream [184, 18]. mainstreamisation [187]. Making [172, 220, 179, 43]. Malaysia [143]. man [151]. manifesto [243]. Mao [142]. Maoist [142]. MapQuest [163]. marginal [184]. Marie [207]. market [103, 109]. markets [106, 75]. Matamoros [232]. Matamoros-Fernández [232]. material [25]. matter [118]. mature [206]. McKelvey [88]. Media [168, 219, 222, 213, 241, 110, 81, 128, 187, 53, 203, 244]. mediated [147]. mediatised [147]. meme [229, 170]. memories [69]. Memory [168, 250]. men [131, 166]. menu [24]. Messaging [232, 220, 85]. Metadata [182]. Methodological [229]. Methods [80, 79, 78, 83]. Michel [69]. mid [72]. mid-1990s [72]. Middle [11]. millennium [237]. Milligan [90]. Ministerial [197]. Ministry [46]. Minitel [121, 60]. Minneapolis [88]. Minnesota [88]. misremediation [98]. missing [6]. Mobile [106]. model [11]. modelled [242]. Modem [219]. modems [66]. Moderation [178, 227, 185]. moderator [138]. modernity [22]. moment [190]. mommy [238]. Moon [210]. Mormon [238]. Movement [42, 133, 144, 55, 138]. Movies [248]. Mullaney [225]. multi [204]. multi-level [204]. multidisciplinary [109]. multilingualism [13]. Multipath [248]. museology [211]. Museum [217, 207, 218, 208, 210]. Museums [215, 213, 212, 206]. my [242]. Myspace [157]. Myth [123].

narratives [97, 201]. Natalie [189]. nation [34]. nationalism [144, 192]. nature [24]. negotiating [186]. neoliberal [196]. ’net [240, 6]. Netherlands [76]. netizens [145, 188, 188]. Network [240, 107, 123, 72, 105, 37, 223, 34, 86]. networked [180, 70]. networking [44, 143]. Networks [132, 194, 104, 120, 105, 96, 151, 52, 35]. neutrality [107, 37]. newsgroup [166]. Nicholas [28]. Niels [90]. nonbinary [241]. nonlinear [25]. nostalgic [157]. Notes [15]. noun [23]. novelty [38]. null [15]. NYU [231].

objectives [173]. offline [223]. Old [202]. Olympics [149]. once [40]. one [232]. one-to-one [232]. ones [171]. Online [183, 125, 41, 146, 198, 209, 46, 223, 208, 21]. only [209]. onset [227]. open [212, 104]. openness [211]. operationalising [12]. organizations [222]. origins [119, 42]. other [105]. Our [61]. Outlining [116]. overseas [149]. own [27]. Oxford [230].

packet [65]. pandemic [208]. Paolo [123]. paperback [214]. paradoxes [106]. part [223]. passwords [161]. Past [168, 42, 221]. Patterns [247]. Paul [89, 27]. people [130]. Performing [80]. periphery [148]. Perl [54]. Perry [172]. persistences [204]. perspective [222, 109]. Peter [70]. Peters [34]. photos [43]. pieces [81]. Pierre [57]. Pierrot [210]. pilot [12]. pioneer [27, 44, 68]. pioneers [238]. piracy [247]. place [112]. Platform [60, 84, 156, 159, 73, 167, 232]. platform-as-infrastructure [73]. platforms [155, 154]. PLATO [7]. play [191]. playful [150]. point [121]. policy [121, 107, 37]. political [175, 82, 42]. politics [130, 240, 165, 117, 96, 237]. Polity [189, 232, 174]. poor [151]. Popeye [248]. popular [144]. Porn [159]. Possessed [88]. Post [137, 193, 19, 142, 22]. post-digital [19]. Post-IPO [137]. post-Mao [142]. post-secular [22]. post-socialist [193]. postmen [65]. potatoes [65]. Powell.pps [249]. Power [111, 132, 35]. pp [231, 189, 225, 224, 214, 230, 88]. practices [227, 221, 143]. Practicing [105]. praxis [142]. PRC [142]. Pre [219, 7]. Pre-History [219]. pre-Internet [7]. precarity [154]. Precorporation [10]. precursors [146]. prejudice [148]. Preserving [207]. Press [231, 60, 169, 89, 225, 88, 232, 224, 214, 230, 189, 174]. pride [148]. primary [39, 134, 249]. Prime [197]. Pristina [210]. Privacy [251, 105]. Privatization [183]. process [8, 107, 178]. Production [168, 41]. Professionals [95]. profiles [173, 30]. project [48, 52]. projects [72]. promise [175, 130]. promises [211]. propaganda [194]. proper [23]. protest [35]. protests [149]. protocol [85]. provider [76]. pseudonymity [30]. public [117, 150, 152]. publics [150]. pure [144]. purges [159].

quarantines [165]. queer [240, 159, 243]. quotation [98].

R [174, 165]. r/WatchRedditDie [165]. racial [119]. Radical [142, 138]. RangKoM [143]. Ranking [168]. Re [2, 233, 157]. reactions [56]. reading [249]. Reagle [169]. real [151]. Realm [50]. rebirths [159]. rec.arts.anime [177]. reconstructing [134, 242, 221]. reconstruction [204]. recounting [81]. Recovering [98]. recovery [216]. Recuber [231]. reddit [165, 179]. Reding [110]. reimagination [216]. related [187]. relay [149]. Relentless [145]. religion [126, 22]. religious [56]. Remembering [188]. Rennie [174]. Renren [161]. representation [242]. representing [171]. research [21, 19, 229, 134, 102, 135]. Researcher [226, 113, 176]. Researching [245].

Resistance [101, 195]. Responsibility [95]. responsible [95]. Rethinking [211, 96]. retrospective [84]. reveries [33]. Review [21, 50, 49, 231, 60, 91, 189, 28, 169, 89, 70, 225, 88, 123, 90, 232, 174, 33, 224, 35, 214, 22, 34, 230, 212, 119]. reviewers [36, 63, 93, 124, 153]. revisionism [3]. revisiting [149, 162]. Revolution [49, 230, 169]. Revolutions [214]. rewards [137]. rhetorics [246]. Rieder [100]. right [227, 145]. rights [55, 166]. riot [179]. Rise [72, 24, 163]. Rising [48]. risks [137]. Roberts [68]. role [208]. romanticism [146]. Ronda [188]. Roots [86, 67]. round [79, 109]. round-doc [79, 109]. roundtable [245, 155]. Routing [86]. running [85]. Russian [192].

S [225, 122]. S-curve [122]. sacred [22]. SAGE [90]. Sam [22]. Samizdat [192]. Sandvig [21]. saved [160]. scaffolding [241]. scale [83]. scenes [57]. Scheirer [224]. scholarly [180]. Sci [247]. Sci-Hub [247]. science [206]. Sciences [49]. screencast [17]. scripts [13]. search [24]. Searching [6]. second [176]. secondary [75]. secrets [21]. secular [22]. Secure [220]. seeing [205]. seeking [147, 167]. selection [76]. selective [118]. self [180]. self-archiving [180]. Selves [70]. Semantics [82]. semiotic [43]. sense [172]. Serres [67]. service [248]. servitude [166]. SGML [98]. Shaping [197]. shareholder [137]. sharing [41, 43, 28]. Sheets [186]. Short [230, 133]. SHOT [245]. should [16]. Signal [85]. significance [160]. silence [239]. Siles [70]. Silicon [228]. Simulmatics [139]. Singapore [201, 202]. site [29]. situated [38]. Situating [193, 154]. skeptical [246]. slash [237]. Social [168, 95, 219, 213, 241, 211, 7, 223]. socialist [193]. society [110, 114, 86]. Socio [75, 47]. Socio-technical [75, 47]. sociotechnical [235]. Socrate} [72]. software [182]. something [191]. sonic [128]. Sorting [168, 205]. sources [39, 134, 249]. Soviet [34]. spaces [204]. Speech [138, 121]. speeches [197]. sphere [201, 150]. Spheres [22]. spiral [239]. stake [184]. standardisation [85]. Standardising [85]. standards [98]. Stanford [224]. Star [42]. state [144]. States [70, 227]. STEM [194]. stopped [24]. Stories [169, 142]. story [77]. strategies [105]. Streaming [239, 74, 248]. structuralism [66]. structure [9]. structures [105]. struggles [29]. studies [83, 135, 229, 60]. studio [248]. Study [80, 38, 43, 12]. studying [21]. Style [186]. subreddits [178]. Surveillance [111]. sustaining [209]. swarms [25]. Swartz [196]. Sweden [247]. switching [65]. system [136]. systems [38, 228, 94].

talk [164]. Talking [53]. Task [235]. Taxing [108]. Team [160]. teargas [35]. Technical [86, 47, 75]. technician [138]. techniques [100]. technocratic [133]. technological [154]. technologies [126, 83, 186, 22]. Technologists [250]. Technology [175, 56, 206, 43, 54, 208]. Tell [20, 10]. Téramond [32]. text [98]. Thanks [36, 63, 93, 124, 153]. therefore [138]. They’re [162]. Things [224]. thinking [228]. third [226]. Thomas [225, 174]. threads [191]. tie [198]. Tiidenberg [189]. timesharing [38]. Timothy [231]. together [248]. Tom [157]. Tomsk [97]. tool [222]. tools [128, 43]. Tor [251, 190, 87]. torch [149]. towns [202]. toxic [185]. Tracing [117, 46, 181]. tracking [83, 135]. tragedy [40]. Trajectories [70, 100]. transformations [75]. Transformed [61]. transforming [102]. Transgender [214]. transition [237]. transnational [52]. treated [16]. trials [190]. Trolls [187]. Tufekci [35]. Tumblr [160, 189]. turn [196, 237]. turning [121, 51]. turning-point [121]. twitch [239].

Twitter [35]. Two [214, 201].

U.S [169]. UK [46]. unclaimed [98]. uncovering [7]. uneasy [34]. unexplored [12]. Union [107]. United [70, 227]. University [88, 224, 214, 230]. Unix [151]. untold [77]. updates [221]. uptake [120]. URLs [205]. useful [39]. Usenet [188, 80]. user [136, 236]. usernames [30]. users [106]. Using [80]. utopianism [95].

vaccine [246]. Valley [228]. value [137, 128, 18]. VarageSale [75]. Vegas [44]. VidAngel [126]. Video [61]. videogames [248]. videos [43]. view [8]. views [197]. vigilantism [147]. Village [162]. villains [145]. Virtual [231, 206]. visibility [239, 152]. visions [143]. Viviane [110]. vocabulary [187].

Wall [54]. Walter [224]. war [194]. wars [39]. was [223]. WatchRedditDie [165]. ways [54]. Weaving [87]. Web [91, 98, 222, 84, 187, 181, 59, 48, 204, 205, 158, 92, 157, 83, 160, 249, 221, 90, 170, 102, 122, 46, 54, 76, 243, 186, 18, 245, 246, 200, 251, 250, 134, 167, 17, 51, 87, 76, 62, 215]. Web-based [134]. websites [218]. Weibo [117]. Welcome [60, 121]. Werner [33]. WhatsApp [232]. where [2]. who [77]. whole [228]. whole-systems [228]. Wi [174]. Wi-Fi [174]. Wide [200, 62]. Wikidata [242]. Wikipedia [190, 203, 50, 169]. Wikipedia.org [203]. Wikipedian [235]. Wilken [174]. Williams [207]. Willred [210]. winning [44]. Wired [118]. within [135]. woman [234]. Women [41, 77, 131, 235, 207, 177, 234, 239]. women-exclusionary [177]. words [27]. Work [50, 134]. World [61, 219, 95, 131, 197, 200, 62, 33]. write [5]. writing [233].

Xena [248]. yak [164]. Yakety [164]. Years [32, 191, 62]. Yelp [162]. York [70, 214]. Yugoslavia [194].

Zen [186]. zeros [171]. Zeynep [35].

References

Brugger:2017:IIH

[1] Niels Brügger, Gerard Goggin, Ian Milligan, and Valérie Schafer. Introduction: Internet histories. Internet Histories, 1(1–2):1–7, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                             tandfonline.com/doi/full/10.1080/
                                             24701475.2017.1317128.

Abbate:2017:WWI

[2] Janet Abbate. What and where is the Internet? (re)defining Internet histories. Internet Histories, 1(1–2):8–14, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                             tandfonline.com/doi/full/10.1080/
                                             24701475.2017.1305836.

Russell:2017:HRB

[3] Andrew L. Russell. Hagiography, revisionism & blasphemy in Internet histories. Internet Histories, 1 (1–2):15–25, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                             tandfonline.com/doi/full/10.1080/
                                             24701475.2017.1298229.

Weber:2017:CL

[4] Marc Weber. A common language. Internet Histories, 1(1–2):26–38, 2017. CODEN ???? ISSN 2470-1483.

                                             URL http://www.tandfonline.com/
                                             doi/full/10.1080/24701475.2017.
                                             1317118.

Turner:2017:CWW

[5] Fred Turner. Can we write a cultural history of the Internet? If so, how? Internet Histories, 1 (1–2):39–46, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1307540.

Driscoll:2017:SMN

[6] Kevin Driscoll and Camille Paloque-Berges. Searching for missing “net histories”. Internet Histories, 1 (1–2):47–59, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1307541.

Jones:2017:PCU

[7] Steve Jones and Guillaume Latzko-Toth. Out from the PLATO cave: uncovering the pre-Internet history of social computing. Internet Histories, 1(1–2):60–69, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1307544.

Braman:2017:IHV

[8] Sandra Braman. Internet Histories: the view from the design process. Internet Histories, 1(1–2):70–78, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1305716.

Streeter:2017:ISF

[9] Thomas Streeter. The Internet as a structure of feeling: 1992–1996. Internet Histories, 1(1–2):79–89, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1306963.

Elmer:2017:PWF

[10] Greg Elmer. Precorporation: or what financialisation can tell us about the histories of the Internet. Internet Histories, 1(1–2):90–96, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1308197.

Allagui:2017:IME

[11] Ilhem Allagui. Internet in the Middle East: an asymmetrical model of development. Internet Histories, 1 (1–2):97–105, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1305715.

Reisdorf:2017:UHO

[12] Bianca C. Reisdorf, William H. Dutton, Whisnu Triwibowo, and Michael E. Nelson. The unexplored history of operationalising digital divides: a pilot study. Internet Histories, 1 (1–2):106–118, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1311165.

McLelland:2017:ECM

[13] Mark McLelland. Early challenges to multilingualism on the Internet: the case of Han characterbased scripts. Internet Histories, 1 (1–2):119–128, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1280889.

Wasserman:2017:AHI

[14] Herman Wasserman. African histories of the Internet. Internet Histories, 1(1–2):129–137, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1308198.

Brunton:2017:NDN

[15] Finn Brunton. Notes from /dev/null. Internet Histories, 1(1–2):138–145, 2017. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2017.
1307059.

Alberts:2017:AAD

[16] Gerard Alberts, Marc Went, and Robert Jansma. Archaeology of the Amsterdam digital city; why digital data are dynamic and should be treated accordingly. Internet Histories, 1 (1–2):146–159, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1309852.

Rogers:2017:DWH

[17] Richard Rogers. Doing Web history with the Internet Archive: screencast documentaries. Internet Histories, 1 (1–2):160–172, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1307542.

Winters:2017:BMD

[18] Jane Winters. Breaking in to the mainstream: demonstrating the value of Internet (and web) histories. Internet Histories, 1(1–2):173–179, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1305713.

Fortunati:2017:DPD

[19] Leopoldina Fortunati. For a dynamic and post-digital history of the Internet: a research agenda. Internet Histories, 1(1–2):180–187, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1305714.

Schafer:2017:TUA

[20] Valérie Schafer. Tell us about . . .. Internet Histories, 1(1–2):188–196, 2017. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2017.
1301132.

Acker:2017:BRD

[21] Amelia Acker. Book review: Digital research confidential: the secrets of studying behavior online, by Hargittai, E. and Sandvig, C. Internet Histories, 1(1–2):197–199, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1313928.

Webster:2017:BRT

[22] Peter Webster. Book review: Technologies of religion. Spheres of the sacred in a post-secular modernity, by Sam Han. Internet Histories, 1 (1–2):200–201, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1316535.

Bay:2017:WIC

[23] Morten Bay. What is “Internet”? The case for the proper noun and why it is important. Internet Histories, 1(3):203–218, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1339860.

Kerssens:2017:WSE

[24] Niels Kerssens. When search engines stopped being human: menu interfaces and the rise of the ideological nature of algorithmic search. Internet Histories, 1(3):219–237, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1337666.

Micali:2017:TNM

[25] Alberto Micali. Towards a nonlinear, material history of digital swarms. Internet Histories, 1(3):238–257, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2017.1344051.

LeGallic:2017:IFC

[26] Stéphanie Le Gallic. Infonie: a forced conversion to the Internet (1995–2000). Internet Histories, 1 (3):258–272, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1344417.

Bay:2017:CPP

[27] Morten Bay. Conversations with a pioneer: Paul Baran in his own words. Internet Histories, 1(3):273–282, 2017. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2017.
1345469.

Halavais:2017:BRA

[28] Alexander Halavais. Book review: The age of sharing, by Nicholas A. John. Internet Histories, 1 (3):283–284, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1346223.

Hellegren:2017:HCD

[29] Z. Isadora Hellegren. A history of crypto-discourse: encryption as a site of struggles to define Internet freedom. Internet Histories, 1 (4):285–311, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1387466.

vanderNagel:2017:UPD

[30] Emily van der Nagel. From usernames to profiles: the development of pseudonymity in Internet communication. Internet Histories, 1 (4):312–331, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1389548.

Brown:2017:LEH

[31] Katherine Brown. LCHC: the evolution of a hybrid collaboratory. Internet Histories, 1(4):332–348, 2017. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2017.1384111.

Siles:2017:YIC

[32] Ignacio Siles. 25 years of the Internet in Central America: an interview with Guy de Téramond. Internet Histories, 1(4):349–358, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2017.1383733.

Rouyan:2017:FRB

[33] Anahita Rouyan. Film review: Lo and behold: reveries of the connected world, directed by Werner Herzog. Internet Histories, 1(4):359–360, 2017. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2017.1382827.

West:2017:BRH

[34] Diana Kurkovsky West. Book review: How not to network a nation: the uneasy history of the Soviet Internet, by Benjamin Peters. Internet Histories, 1(4):361–363, 2017. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2017.1382828.

Summers:2017:BRT

[35] Ed Summers. Book review: Twitter and teargas: the power and fragility of networks protest, by Zeynep Tufekci. Internet Histories, 1(4):364–367, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2017.1383639.

Anonymous:2017:TR

[36] Anonymous. Thanks to reviewers. Internet Histories, 1(4):368–369, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2017.1394109.

Mailland:2018:BIP

[37] Julien Mailland. Building Internet policy on history: lessons of the forgotten 1981 network neutrality debate. Internet Histories, 2(1–2):1–19, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1450702.

DePaoli:2018:ECN

[38] Stefano De Paoli. The engineer– criminologist and “the novelty of cybercrime”: a situated genealogical study of timesharing systems. Internet Histories, 2(1–2):20–37, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1448498.

Hocquet:2018:MLA

[39] Alexandre Hocquet and Frédéric Wieber. Mailing list archives as useful primary sources for historians: looking for flame wars. Internet Histories, 2(1–2):38–54, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1456741.

Dourish:2018:OFI

[40] Paul Dourish. The once and future Internet: infrastructural tragedy and ambiguity in the case of IPv6. Internet Histories, 2(1–2):55–74, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1455311.

Fortunati:2018:WKC

[41] Leopoldina Fortunati. Women’s knowledge co-production and sharing in online communities. Internet Histories, 2(1–2):75–97, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1459132.

Musso:2018:OPD

[42] Marta Musso and Marzia Maccaferri. At the origins of the political discourse of the 5-Star Movement (M5S): Internet, direct democracy and the “future of the past”. Internet Histories, 2(1–2):98–120, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1457295.

Poulsen:2018:BST

[43] Søren Vigild Poulsen. Becoming a semiotic technology — a historical study of Instagram’s tools for making and sharing photos and videos. Internet Histories, 2(1–2):121–139, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1459350.

Bay:2018:CPL

[44] Morten Bay. Conversation with a pioneer: Leonard Kleinrock on the early days of networking, the ARPANET … and winning in Las Vegas. Internet Histories, 2(1–2):140–152, 2018. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2018.1446239.

Winters:2018:E

[45] Jane Winters. Editorial. Internet Histories, 2(1–2):153–155, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1459080.

Raffal:2018:TOD

[46] Harry Raffal. Tracing the online development of the Ministry of Defence and Armed Forces through the UK web archive. Internet Histories, 2 (1–2):156–178, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1456739.

Ben-David:2018:IAS

[47] Anat Ben-David and Adam Amram. The Internet Archive and the socio-technical construction of historical facts. Internet Histories, 2 (1–2):179–201, 2018. CODEN ????

ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2018.1455412.

Kunze:2018:CCE

[48] Svenja Kunze and Brendan Power. Capturing commemoration: the 1916 Easter Rising web archive project. Internet Histories, 2(1–2):202–217, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1446238.

Cowls:2017:BRI

[49] Josh Cowls and Evan Higgins. Book review: The Internet Revolution in the Sciences and Humanities, by Alan G. Gross and Joseph E. Harmon. Internet Histories, 2(1–2):218–220, 2017. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2017.1419841.

Chakraborty:2017:RWW

[50] Anwesha Chakraborty. Review of Wikipedia, Work and Capitalism: A Realm of Freedom?, by Arwid Lund. Internet Histories, 2(1–2):221–223, 2017. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2017.
                                        1415279.

Schafer:2018:TDI

[51] Valérie Schafer and Benjamin G. Thierry. The 90s as a turning decade for Internet and the Web. Internet Histories, 2(3–4):225–229, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1521060.

Siles:2018:ITP

[52] Ignacio Siles. The Internet as a transnational project: connecting Central America through computer networks (1990–1996). Internet Histories, 2(3–4):230–246, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1500793.

Horbinski:2018:TLH

[53] Andrea Horbinski. Talking by letter: the hidden history of female media fans on the 1990s Internet. Internet Histories, 2(3–4):247–263, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1500794.

Stevenson:2018:HIB

[54] Michael Stevenson. Having it both ways: Larry Wall, Perl and the technology and culture of the early web. Internet Histories, 2(3–4):264–280, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1495810.

Petin:2018:BDA

[55] Patrick Pétin and Félix Tréguer. Building and defending the alternative Internet: the birth of the digital rights movement in France. Internet Histories, 2(3–4):281–298, 2018. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1521059.

Webster:2018:TER

[56] Peter Webster. Technology, ethics and religious language: early Anglophone Christian reactions to “cyberspace”. Internet Histories, 2(3– 4):299–314, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2018.1468976.

Paloque-Berges:2018:BSI

[57] Camille Paloque-Berges, Victoria Peuvrelle, and Valérie Schafer. Behind the scenes: an interview with Pierre Beyssac. Internet Histories, 2 (3–4):315–327, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2018.1431496.

Schafer:2018:ICD

[58] Valérie Schafer. Interfacing counterculture and digital cultures: an interview with Geert Lovink. Internet Histories, 2(3–4):328–338, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1526525.

Hockx-Yu:2018:WH

[59] Helen Hockx-Yu. The web as history. Internet Histories, 2(3–4):340–342, 2018. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2018.
                                        1509579.

Gebeil:2018:BRJ

[60] Sophie Gebeil. Book review: Julien Mailland & Kevin Driscoll, Minitel Welcome to the Internet, Platform Studies, MIT Press, 2017. Internet Histories, 2(3–4):342–345, 2018. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2018.1496310.

Bonfiglioli:2018:BBH

[61] Rudi Bonfiglioli. Bit by bit: How video games transformed our world by Andrew ervin. Internet Histories, 2 (3–4):346–349, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1449621.

Winters:2018:WHF

[62] Jane Winters. Web 25: histories from the first 25 years of the World Wide Web. Internet Histories, 2 (3–4):350–352, 2018. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1439149.

Anonymous:2018:TR

[63] Anonymous. Thanks to reviewers. Internet Histories, 2(3–4):353–354, 2018. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2018.
1535824.

Paloque-Berges:2019:A

[64] Camille Paloque-Bergès and Valérie Schafer. Arpanet (1969–2019). Internet Histories, 3(1):1–14, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1560921.

Bay:2019:HPP

[65] Morten Bay. Hot potatoes and postmen: how packet switching became ARPANET’s greatest legacy. Internet Histories, 3(1):15–30, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1544726.

McKelvey:2019:ABD

[66] Fenwick McKelvey and Kevin Driscoll. ARPANET and its boundary devices: modems, IMPs, and the interstructuralism of infrastructures. Internet Histories, 3(1):31–50, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1548138.

Schafer:2019:BRA

[67] Valérie Schafer, Sarah Cooper, and Camille Paloque-Bergès. Back to the roots of ARPANET and Internet history with Alexandre Serres. Internet Histories, 3(1):51–67, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2018.1560680.

Bay:2019:CPL

[68] Morten Bay. Conversation with a pioneer: Larry Roberts on how he led the design and construction of the ARPANET. Internet Histories, 3(1):68–80, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1544727.

Paloque-Berges:2019:FMA

[69] Camille Paloque-Bergès and Valérie Schafer. French memories about the ARPANET: a conversation with Michel Élie and Gérard Le Lann. Internet Histories, 3(1):81–97, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2018.1536934.

Locatelli:2019:BRN

[70] Elisabetta Locatelli. Book review: Networked Selves: Trajectories of Blogging in the United States and France, by I. Siles, New York, Peter Lang, 2017. ISBN-978-1-4331-4710-4. Internet Histories, 3(1):98–100, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1565555.

Lauscher:2019:LBH

[71] Anne Lauscher. Life 3.0: being human in the age of artificial intelligence. Internet Histories, 3(1):101–103, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1565556.

Bory:2019:INH

[72] Paolo Bory. The Italian network hopes: Rise and fall of the Socrate and Iperbole projects in the mid-1990s. Internet Histories, 3(2):105–122, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1596407.

Helmond:2019:FED

[73] Anne Helmond, David B. Nieborg, and Fernando N. van der Vlist. Facebook’s evolution: development of a platform-as-infrastructure. Internet Histories, 3(2):123–146, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1593667.

Burroughs:2019:CLS

[74] Benjamin Burroughs. A cultural lineage of streaming. Internet Histories, 3(2):147–161, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1576425.

Lingel:2019:STT

[75] Jessa Lingel. Socio-technical transformations in secondary markets: a comparison of craigslist and Varage-Sale. Internet Histories, 3(2):162–179, 2019. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2018.
                                        1478267.

Teszelszky:2019:WAN

[76] Kees Teszelszky. Web archaeology in The Netherlands: the selection and harvest of the Dutch web incunables of provider Euronet (1994–2000). Internet Histories, 3(2):180–194, 2019. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2019.1603951.

Hockx-Yu:2019:BBU

[77] Helen Hockx-Yu. Broad band, the untold story of the women who made the Internet. Internet Histories, 3 (2):195–198, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1598078.

Brugger:2019:IIH

[78] Niels Brügger and Ian Milligan. Introduction: Internet histories and computational methods. Internet Histories, 3(3–4):199–201, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1675405.

Brugger:2019:IHC

[79] Niels Brügger, Ian Milligan, Anat Ben-David, Sophie Gebeil, Federico Nanni, Richard Rogers, William J.

Turkel, Matthew S. Weber, and Peter Webster. Internet histories and computational methods: a “rounddoc” discussion. Internet Histories, 3(3–4):202–222, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1639352.

Dame-Griff:2019:HPE

[80] Avery Dame-Griff. Herding the ‘performing elephants:’ using computational methods to study Usenet. Internet Histories, 3(3–4):223–244, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1652456.

Dieter:2019:CPR

[81] Michael Dieter, David Gauthier, and Marc Tuters. Conversation pieces: On recounting new media art mailinglist cultures. Internet Histories, 3 (3–4):245–274, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1674580.

Hosl:2019:SIP

[82] Maximilian Hösl. Semantics of the Internet: a political history. Internet Histories, 3(3–4):275–292, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1656921.

Nielsen:2019:ECM

[83] Janne Nielsen. Experimenting with computational methods for large-scale studies of tracking technologies in web archives. Internet Histories, 3 (3–4):293–315, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1671074.

Ben-David:2019:FCP

[84] Anat Ben-David. 2014 not found: a cross-platform approach to retrospective web archiving. Internet Histories, 3(3–4):316–342, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1654290.

Ermoshina:2019:SRC

[85] Ksenia Ermoshina and Francesca Musiani. “Standardising by running code”: the Signal protocol and de facto standardisation in end-to-end encrypted messaging. Internet Histories, 3(3–4):343–363, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1654697.

Fidler:2019:EIR

[86] Bradley Fidler. The evolution of Internet routing: Technical roots of the Network Society. Internet Histories, 3(3–4):364–387, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1661583.

Schriner:2019:WDW

[87] John Schriner. Weaving the Dark Web: Legitimacy on Freenet, Tor, and I2P. Internet Histories, 3(3– 4):388–390, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1623002.

Musiani:2019:BRF

[88] Francesca Musiani. Book review: Fenwick McKelvey. 2018, Internet Daemons. Digital Communications Possessed, Minneapolis: University of Minnesota Press, 320 pp. Internet Histories, 3(3–4):391–393, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1611241.

Lecuyer:2019:BRG

[89] Christophe Lécuyer. Book review: GPS, by Paul Ceruzzi, MIT Press, Cambridge, MA., 2018. Internet Histories, 3(3–4):394–395, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1629192.

Pedersen:2019:BRS

[90] Max Odsbjerg Pedersen and Helle Strandgaard Jensen. Book review: The SAGE handbook of web history, edited by Niels Brügger and Ian Milligan. Internet Histories, 3(3–4):396–399, 2019. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2019.1657643.

Gebeil:2019:BRC

[91] Sophie Gebeil. Book review: En construction, la Fabrique française d’Internet et du Web dans les années 1990, Ina, coll. Études & Controverses. Internet Histories, 3(3– 4):400–403, 2019. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2019.1672952.

Millward:2019:AWD

[92] Gareth Millward. The archived web: doing history in the digital age. Internet Histories, 3(3– 4):404–406, 2019. CODEN ????

ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1673058.

Anonymous:2019:TR

[93] Anonymous. Thanks to reviewers. Internet Histories, 3(3–4):407–408, 2019. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2019.
                                        1677399.

Stevenson:2020:LSI

[94] Michael Stevenson and Anne Helmond. Legacy systems: Internet histories of the abandoned, discontinued and forgotten. Internet Histories, 4(1):1–5, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1725854.

Finn:2020:CWD

[95] Megan Finn and Quinn DuPont. From closed world discourse to digital utopianism: the changing face of responsible computing at Computer Professionals for Social Responsibility (1981–1992). Internet Histories, 4(1):6–31, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1725851.

Kerssens:2020:RLI

[96] Niels Kerssens. Rethinking legacies in Internet history: Euronet, lost (inter)networks, EU politics. Internet Histories, 4(1):32–48, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1701919.

Kolozaridi:2020:NWI

[97] Polina Kolozaridi and Dmitry Muravyov. The narratives we inherit: the local and global in Tomsk’s Internet history. Internet Histories, 4(1):49–65, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1723980.

Ammann:2020:RWU

[98] Rudolf Ammann. Recovering the web’s unclaimed legacy of academic text standards: SGML, HTML, and the misremediation of quotation. Internet Histories, 4(1):66–86, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1725853.

Jones:2020:CLC

[99] Meg Leta Jones. Cookies: a legacy of controversy. Internet Histories, 4(1):87–104, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1725852.

Stevenson:2020:HTA

[100] Michael Stevenson and Anne Helmond.

The historical trajectories of algorithmic techniques: an interview with Bernhard Rieder. Internet Histories, 4(1):105–114, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1723345.

Byrne:2020:DFC

[101] Helena Byrne. Digital football cultures:

Fandom, identities and resistance. Internet Histories, 4(1):115–116, 2020. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2019.1678910.

PhD:2020:HAA

[102] Max Kemman PhD. History in the age of abundance? How the web is transforming historical research. Internet Histories, 4(1):117–119, 2020. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2019.1681715.

Fickers:2020:EIE

[103] Andreas Fickers and Valérie Schafer. Editorial: the Internet and the EU market. Internet Histories, 4 (2):121–124, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1749808.

Henrich-Franke:2020:ECL

[104] Christian Henrich-Franke. EC competition law and the idea of “open networks” (1950s–1980s). Internet Histories, 4(2):125–141, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1743045.

Jones:2020:PPO

[105] Meg Leta Jones and Kevin Ackermann. Practicing privacy on other networks: network structures, economic arrangements, and identity strategies before cookies. Internet Histories, 4(2):142–160, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1747344.

Goggin:2020:MPE

[106] Gerard Goggin. Mobile paradoxes: European emergence of mobile Internet, users, and markets. Internet Histories, 4(2):161–177, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1741968.

Gadringer:2020:NNE

[107] Stefan Gadringer. Network neutrality in the European Union: a communications policy process analysis. Internet Histories, 4(2):178–195, 2020. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2020.1749807.

Danescu:2020:TIA

[108] Elena Danescu. Taxing intangible assets: issues and challenges for a digital Europe. Internet Histories, 4(2):196–216, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1749806.

Schafer:2020:IEM

[109] Valérie Schafer, Andreas Fickers, David Howarth, Francesca Musiani, Julia Pohle, and Dwayne Winseck. “The Internet and the European market” from a multidisciplinary perspective: a “round-doc” discussion. Internet Histories, 4(2):217–227, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1739360.

Danescu:2020:VRH

[110] Elena Danescu. Viviane Reding on her action in the field of the information society and media (2004–2010). Interview by Elena Danescu. Internet Histories, 4(2):228–241, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1742533.

Sam:2020:ASC

[111] Jillet Sarah Sam. The age of surveillance capitalism: The fight for a human future at the new frontier of power. Internet Histories, 4(2):242–244, 2020. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2019.1706138.

Webster:2020:HPD

[112] Peter Webster. A history of place in the digital age. Internet Histories, 4(2):245, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1731250.

Brugger:2020:IHE

[113] Niels Brügger, Gerard Goggin, Ian Milligan, and Valérie Schafer. Internet Histories early career researcher award. Internet Histories, 4(3):247–251, 2020. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2020.1778309.

Haristya:2020:ECS

[114] Sherly Haristya. The efficacy of civil society in global Internet governance. Internet Histories, 4(3):252–270, 2020. CODEN ???? ISSN 2470-1483.

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/24701475.2020.
                                        1769892.

Grosse:2020:LFC

[115] Meghan Grosse. Laying the foundation for a commercialized Internet: international Internet governance in the 1990s. Internet Histories, 4 (3):271–286, 2020. CODEN ????

ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769890.

Nunes:2020:OHC

[116] Raul Nunes. Outlining the history of cyberactivism in Brazil. Internet Histories, 4(3):287–303, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769893.

Jia:2020:TWC

[117] Lianrui Jia and Xiaofei Han. Tracing Weibo (2009–2019): The commercial dissolution of public communication and changing politics. Internet Histories, 4(3):304–332, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769894.

Ferrari:2020:BMB

[118] Elisabetta Ferrari. Bodies that matter, bodies that don’t: selective disembodiment in the early Wired magazine (1993–1997). Internet Histories, 4(3):333–348, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1769891.

Mari:2020:REE

[119] Will Mari. A review essay: examining the fraught racial, gendered and class-based origins of the early Internet and its antecedents. Internet Histories, 4(3):349–353, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1758436.

Jones:2020:FCU

[120] Steve Jones and David W. Park. The field of communication’s uptake of computers, networks, and the Internet: 1970–2000. Internet Histories, 4(4):355–372, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2019.1704495.

Bellon:2020:GMW

[121] Anne Bellon. “Goodbye Minitel, welcome to the Internet”. The Hourtin speech as a turning-point for French Internet policy. Internet Histories, 4(4):373–389, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1731783.

Priestley:2020:IWE

[122] Maria Priestley, T. J. Sluckin, and Thanassis Tiropanis. Innovation on the web: the end of the S-curve? Internet Histories, 4(4):390–412, 2020. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2020.1747261.

Musiani:2020:BRP

[123] Francesca Musiani. Book review: Paolo Bory. 2020, The Internet Myth: From the Internet Imaginary to Network Ideologies. Internet Histories, 4(4):413–416, 2020. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1806481.

Anonymous:2020:TR

[124] Anonymous. Thanks to reviewers. Internet Histories, 4(4):417–420, 2020. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2020.1840157.

Eriksson:2021:GOC

[125] Maria Eriksson and Guillaume Heuguet.

Genealogies of online content identification — an introduction. Internet Histories, 5(1):1–7, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1878649.

Feller:2021:VCF

[126] Gavin Feller and Andrew Ventimiglia.

VidAngel: Content filtering technologies, religion, and American copyright law. Internet Histories, 5(1):8–29, 2021. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2020.
1831198.

Justie:2021:LHC

[127] Brian Justie. Little history of CAPTCHA. Internet Histories, 5 (1):30–47, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1831197.

Ernst:2021:MEV

[128] Wolfgang Ernst. The media epistemic value of sonic analytics tools. A commentary. Internet Histories, 5(1):48–56, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1862528.

Eriksson:2021:IAK

[129] Maria Eriksson and Guillaume Heuguet.

Interview with Aleksandra Kaminska. Internet Histories, 5(1):57–70, 2021. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2021.
1878650.

Bodker:2021:IPP

[130] Henrik Bødker. An Internet for the people: the politics and promise of craigslist. Internet Histories, 5 (1):71–73, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1827615.

Hockx-Yu:2021:IWD

[131] Helen Hockx-Yu. Invisible women, data bias in a world designed for men. Internet Histories, 5(1):74–78, 2021. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2020.1837580.

Flensburg:2021:NPA

[132] Sofie Flensburg and Signe Sophus Lai. Networks of power. Analysing the evolution of the Danish Internet infrastructure. Internet Histories, 5(2):79–100, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1759010.

Beltramini:2021:ATA

[133] Enrico Beltramini. Against technocratic authoritarianism. A short intellectual history of the cypherpunk movement. Internet Histories, 5 (2):101–118, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1731249. See correc-

tion [140] supplying two additional references.

Hodges:2021:FAE

[134] James A. Hodges. Forensic approaches to evaluating primary sources in Internet history research: reconstructing early Web-based archival work (1989–1996). Internet Histories, 5 (2):119–134, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1784539.

Puschmann:2021:FCA

[135] Cornelius Puschmann and Christian Pentzold. A field comes of age: tracking research on the Internet within communication studies, 1994 to 2018. Internet Histories, 5(2):135–153, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2020.1749805.

Kushner:2021:IUH

[136] Scott Kushner. The instrumentalised user: human, computer, system. Internet Histories, 5(2):154–170, 2021. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2020.1810395.

Bilic:2021:GPI

[137] Paško Bilić and Toni Prug. Google’s post-IPO development: risks, rewards, and shareholder value. Internet Histories, 5(2):171–189, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2020.1864959.

Mommeja:2021:ETR

[138] Julie Momméja. “I am an engineer and therefore a radical”: an interview with Lee Felsenstein, from Free Speech Movement technician to Homebrew Computer Club moderator. Internet Histories, 5(2):190–203, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1917903.

Kotliar:2021:ITH

[139] Dan M. Kotliar. If then: How the Simulmatics Corporation invented the future. Internet Histories, 5 (2):204–206, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1878651.

Anonymous:2021:C

[140] Anonymous. Correction. Internet Histories, 5(2):??, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2020.1860392. See [133].

Goggin:2021:AIH

[141] Gerard Goggin, Haiqing Yu, and Kwang-Suk Lee. Asian Internet histories: an introduction. Internet Histories, 5(3–4):207–213, 2021. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2021.1988239.

Qiu:2021:RPC

[142] Jack Linchuan Qiu and Hongzhe Wang. Radical praxis of computing in the PRC: forgotten stories from the Maoist to post-Mao era. Internet Histories, 5 (3–4):214–229, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1949817.

Stevens:2021:RJI

[143] Hallam Stevens. From RangKoM and JARING to the Internet: visions and practices of electronic networking in Malaysia, 1983–1996. Internet Histories, 5(3–4):230–247, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1920126.

Lee:2021:BDS

[144] Kwang-Suk Lee. Between the developmental state and popular nationalism: the pure Hangul movement in the early history of the Korean Internet. Internet Histories, 5 (3–4):248–267, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1951960.

Lee:2021:RVF

[145] Hojeong Lee. Relentless villains or fervent netizens?: The alt-right community in Korea, Ilbe. Internet Histories, 5(3–4):268–286, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.1953756.

Fujioka:2021:DCR

[146] Brett J. Fujioka and Julia R. DeCook.

Digital cynical romanticism: Japan’s 2channel and the precursors to online extremist cultures. Internet Histories, 5(3–4):287–303, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1919966.

Huang:2021:MMJ

[147] Qian Huang. The mediated and mediatised justice-seeking: Chinese digital vigilantism from 2006 to 2018. Internet Histories, 5(3–4):304–322, 2021. CO-DEN ???? ISSN 2470-1483. URL http:

//www.tandfonline.com/doi/full/
10.1080/24701475.2021.1919965.

Parray:2021:CPP

[148] M. Imran Parray. Choking the ‘periphery’: pride and prejudice in India’s globalizing Internet imaginary. Internet Histories, 5(3–4):323–340, 2021. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2021.1984732.

Guo:2021:DAA

[149] Shaohua Guo. Discursive activism in the age of BBS: revisiting overseas Chinese protests during the 2008 Beijing Olympics torch relay. Internet Histories, 5(3–4):341–358, 2021. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2021.1951962.

Xie:2021:PSM

[150] Emilie Xie, Maxwell Foxman, and Shuo Xu. From public sphere to magic circle: playful publics on the Chinese Internet. Internet Histories, 5 (3–4):359–375, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1982166.

Paloque-Berges:2021:RPM

[151] Camille Paloque-Bergès. The real “poor man’s Arpanet”? A conversation about Unix networks with Kilnam Chon, godfather of the Asian Internet. Internet Histories, 5(3– 4):376–388, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1982167.

Yu:2021:ECI

[152] Haiqing Yu. The evolution of the Chinese Internet: Creative visibility in the digital public. Internet Histories, 5 (3–4):389–391, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1959213.

Anonymous:2021:TR

[153] Anonymous. Thanks to reviewers. Internet Histories, 5(3–4):392–394, 2021. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2021.
2008676.

McCammon:2022:SDD

[154] Muira McCammon and Jessa Lingel.

Situating dead-and-dying platforms: technological failure, infrastructural precarity, and digital decline. Internet Histories, 6(1–2):1–13, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2071395.

McCammon:2022:DDP

[155] Muira McCammon, Diami Virgilio, Cody Ogden, Kevin Ackermann, Ethan Zuckerman, Robert Gehl, Saima Akhtar, Sultan Al-Azri, Catherine Knight Steele, Amber M. Hamilton, Anat Ben-David, Sarah Wasserman, Sara Namusoga-Kaale, and Joy Lisi Rankin. Dead-and-dying platforms: a roundtable. Internet Histories, 6(1–2):14–30, 2022. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2071396.

Corry:2021:WDP

[156] Frances Corry. Why does a platform die? Diagnosing platform death at Friendster’s end. Internet Histories, 6(1–2):31–47, 2021. CODEN ????

ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1985360.

Miltner:2021:TUA

[157] Kate M. Miltner and Ysabel Gerrard. “Tom had us all doing frontend web development”: a nostalgic (re)imagining of Myspace. Internet Histories, 6(1–2):48–67, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1985836.

Mikolajewska-Zajac:2022:FDC

[158] Karolina Mikołajewska-Zaja̧c and Attila Márton. The four deaths of Couchsurfing and the changing ecology of the web. Internet Histories, 6(1–2):68–89, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2057751.

Floegel:2021:PBP

[159] Diana Floegel. Porn bans, purges, and rebirths: the biopolitics of platform death in queer fandoms. Internet Histories, 6(1–2):90–112, 2021. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2021.1985833.

Ogden:2021:EIC

[160] Jessica Ogden. “Everything on the Internet can be saved”: Archive Team, Tumblr and the cultural significance of web archiving. Internet Histories, 6(1–2):113–132, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1985835.

Jia:2022:FPL

[161] Lianrui Jia. Forgotten passwords and Long-Gone exes: the life and death of Renren. Internet Histories, 6 (1–2):133–153, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2065153.

Kneese:2022:TDY

[162] Tamara Kneese. “They’re describing Yelp in 1992!”: revisiting the Blacksburg Electronic Village. Internet Histories, 6(1–2):154–171, 2022. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2049069.

Wilken:2022:RFM

[163] Rowan Wilken. The rise and fall of MapQuest. Internet Histories, 6 (1–2):172–190, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2057752.

Montalbano:2021:YYD

[164] Kathryn Montalbano. “Yakety yak:

Don’t talk back”: An autopsy of anonymity gone awry. Internet Histories, 6(1–2):191–205, 2021. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2021.2020470.

DeCook:2021:RWP

[165] Julia R. DeCook. r/WatchRedditDie and the politics of reddit’s bans and quarantines. Internet Histories, 6 (1–2):206–222, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1997179.

deConing:2022:LIS

[166] Alexis de Coning. A ‘lifetime of indentured servitude:’ rights, labor, and gender anxieties in a dead men’s rights newsgroup. Internet Histories, 6(1–2):223–236, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2038867.

Mackinnon:2022:DGS

[167] Katie Mackinnon. The death of GeoCities: seeking destruction and platform eulogies in Web archives. Internet Histories, 6(1–2):237–252, 2022. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2051331.

Allmann:2021:SMA

[168] Kira Allmann. Social media and the automatic production of memory: Classification, ranking, and sorting of the past. Internet Histories, 6(1–2):253–256, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1953885.

Hockx-Yu:2021:BRW

[169] Helen Hockx-Yu. Book review: Wikipedia @ 20, Stories of an Incomplete Revolution, edited by Joseph Reagle and Jackie Koerner, The MIT Press, Cambridge, MA, 2020. U.S. $27.95. Internet Histories, 6 (1–2):257–261, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1932273.

Pettis:2022:KYM

[170] Ben T. Pettis. Know Your Meme and the homogenization of web history. Internet Histories, 6(3):263–279, 2022. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2021.
1968657.

Ramati:2022:ABD

[171] Ido Ramati. Aleph-bet, dits-and-dahs, zeros and ones: representing Hebrew in character code. Internet Histories, 6(3):280–297, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1953757.

Buozis:2022:MCS

[172] Michael Buozis. Making common sense of cyberlibertarian ideology: The journalistic consecration of John Perry Barlow. Internet Histories, 6 (3):298–314, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1943994.

Jarvis:2022:CIO

[173] Craig Jarvis. Cypherpunk ideology: objectives, profiles, and influences (1992–1998). Internet Histories, 6(3):315–342, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.1935547.

Rikitianskaia:2022:BRT

[174] Maria Rikitianskaia. Book review:

Thomas, J., Wilken, R., & Rennie, E. (2021). Wi-Fi. Cambridge: Polity press. Internet Histories, 6 (3):343–344, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2031700.

Aidinoff:2022:PAT

[175] Marc Aidinoff. The promise of access: Technology, inequality, and the political economy of hope. Internet Histories, 6(3):345–348, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1936780.

Brugger:2022:IHS

[176] Niels Brügger, Gerard Goggin, Ian Milligan, and Valérie Schafer. Internet histories second early career researcher award. Internet Histories, 6(4):349–351, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2146569.

Petit:2022:DFA

[177] Aurélie Petit. “Do female anime fans exist?” The impact of women-exclusionary discourses on rec.arts.anime. Internet Histories, 6(4):352–368, 2022. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2022.2109265.

deAndrade:2022:MAB

[178] Daniela Linkevicius de Andrade and Demival Vasques Filho. Moderation and authority-building process: the dynamics of knowledge creation on history subreddits. Internet Histories, 6(4):369–390, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2022.2110564.

Harel:2022:AMD

[179] Terri Lee Harel. Archives in the making: documenting the January 6 Capitol riot on Reddit. Internet Histories, 6(4):391–411, 2022. CODEN ????

ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2103989.

MacDonald:2022:INS

[180] Corina MacDonald. Imagining networked scholarly communication: selfarchiving, academic labour, and the early Internet. Internet Histories, 6(4):412–431, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2103987.

Hegarty:2022:IAW

[181] Kieran Hegarty. The invention of the archived web: tracing the influence of library frameworks on web archiving infrastructure. Internet Histories, 6(4):432–451, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2103988.

Jansma:2022:MDD

[182] Robert Jansma. Metadata dating the digital city: a software archaeological approach. Internet Histories, 6(4):452–472, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2109285.

Pettis:2022:NLL

[183] Ben T. Pettis. The new laws of love: Online dating and the privatization of intimacy. Internet Histories, 6(4):473–476, 2022. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2075156.

Clavert:2023:BMM

[184] Frédéric Clavert and Valérie Schafer. Between marginal and mainstream. Communities and ecosystems at stake. Internet Histories, 7(1):1–5, 2023. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2153399.

Passmann:2023:HCT

[185] Johannes Paßmann, Anne Helmond, and Robert Jansma. From healthy communities to toxic debates: Disqus’ changing ideas about comment moderation. Internet Histories, 7(1):6–26, 2023. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2022.2105123.

Wilson:2023:DNA

[186] Derren Wilson, Saeed-Ul Hassan, Naif Radi Aljohani, Anna Visvizi, and Raheel Nawaz. Demonstrating and negotiating the adoption of web design technologies: Cascading Style Sheets and the CSS Zen Garden. Internet Histories, 7(1):27–46, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2055274.

Gomez-Mejia:2023:MWR

[187] Gustavo Gomez-Mejia. The mainstreamisation of web related vocabulary: “Buzz” and “Trolls” in French media archives. Internet Histories, 7(1):47–75, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2089835.

Miller:2023:RNI

[188] Tristan Miller, Camille Paloque-Bergès, and Avery Dame-Griff. Remembering netizens: an interview with Ronda Hauben, co-author of Netizens: on the history and impact of Usenet and the Internet (1997). Internet Histories, 7(1):76–98, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2123120.

Gekker:2023:BRT

[189] Alex Gekker. Book review: Tumblr, By Katrin Tiidenberg, Natalie Ann Hendry, and Crystal Abidin. Polity press: 2021, 240 pp., ISBN 978-1-5095-4108-9. Internet Histories, 7(1):99–103, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2082060.

Gorissen:2023:WWM

[190] Sebastiaan Gorissen and Robert W.

Gehl. When Wikipedia met Tor: trials of legitimacy at a key moment in Internet history. Internet Histories, 7(2):105–121, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2021.2015967.

McKitrick:2023:LPS

[191] Brian McKitrick, Martin Gibbs, Melissa J.

Rogerson, Bjørn Nansen, and Charlotte Pierce. Let’s play something awful: a historical analysis of 14 years of threads. Internet Histories, 7(2):122–140, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2022.2084823.

Mikhailov:2023:EIS

[192] Dmitry Mikhailov, Nikolay Ternov, and Ivan Bobrov. Emigration to the Internet: “Samizdat” and the genesis of contemporary Russian nationalism. Internet Histories, 7(2):141–160, 2023. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2022.2137756.

Bareikyte:2023:SII

[193] Miglė Bareikytė. Situating the Internet as infrastructure: the case of postsocialist Lithuania. Internet Histories, 7(2):161–178, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.2000246.

Brautovic:2023:FPW

[194] Mato Brautović. The first propaganda war through computer networks: STEM academia and the breakup of Yugoslavia. Internet Histories, 7 (2):179–198, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2021.1992992.

Kurzmeier:2023:RCD

[195] Michael Kurzmeier. Resistance to the current: the dialectics of hacking. Internet Histories, 7(2):199–201, 2023. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2023.2188765.

Buozis:2023:IFF

[196] Michael Buozis. Internet freedom, free culture, and free information: Aaron Swartz and cyberlibertarianism’s neoliberal turn. Internet Histories, 7 (3):203–219, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2235131.

Tandon:2023:SDW

[197] Lasya Tandon. Shaping the digital world: views on the Internet in India in Prime Ministerial speeches (1998–2019). Internet Histories, 7 (3):220–236, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2232212.

John:2023:HFO

[198] Nicholas John and Dekel Katz. A history of features for online tie breaking, 1997–2021. Internet Histories, 7(3):237–253, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2214466.

Nabben:2023:CGI

[199] Kelsie Nabben. Cryptoeconomics as governance: an intellectual history from “Crypto Anarchy” to “Cryptoeconomics”. Internet Histories, 7 (3):254–276, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2183643.

Fomasi:2023:GAB

[200] Martin Fomasi, Deborah Barcella, Eleonora Benecchi, and Gabriele Balbi. Genealogy of an archive. The birth, construction, and development of the World Wide Web collection at CERN. Internet Histories, 7(3):277–294, 2023. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2023.2238254.

Lee:2023:BTA

[201] Howard Lee and Terence Lee. Between two acts: competing narratives, activism and governance in Singapore’s digital sphere. Internet Histories, 7(4):295–312, 2023. CODEN ????

ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2232214.

Liew:2023:OCT

[202] Kai Khiun Liew. Old “cowboy towns”: enduring democratic enclaves as Singapore’s alternative digital histories, 1994–2011. Internet Histories, 7 (4):313–332, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2236909.

Jankowski:2023:WIN

[203] Steve Jankowski. The Wikipedia imaginaire: a new media history beyond Wikipedia.org (2001–2022). Internet Histories, 7(4):333–353, 2023. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2023.2246261.

Lobbe:2023:CDW

[204] Quentin Lobbé. Continuity and discontinuity in web archives: a multi-level reconstruction of the firsttuesday community through persistences, continuity spaces and web cernes. Internet Histories, 7(4):354–385, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2254050.

Maemura:2023:SUS

[205] Emily Maemura. Sorting URLs out: seeing the web through infrastructural inversion of archival crawling. Internet Histories, 7(4):386–401, 2023. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2023.2258697.

Bowen:2023:EVS

[206] Jonathan P. Bowen, Ann Borda, Giuliano Gaia, and Stefania Boiano. Early virtual science museums: when the technology is not mature. Internet Histories, 8(1–2):9–33, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2298155.

Kyritsis:2023:PIM

[207] Ismini Kyritsis and Karin de Wild.

Preserving the international museum of women: an interview with Marie Williams Chant. Internet Histories, 8(1–2):34–45, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2267810.

Zhang:2023:ARD

[208] Minying Zhang and Peng Liu. An analysis of the role of digital technology in the online exhibition of the art museum in the Guangdong–Hong Kong– Macao Greater Bay Area during the COVID-19 pandemic. Internet Histories, 8(1–2):62–85, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2286871.

PhD:2023:HFG

[209] Rene Alberto G. Cepeda Ph.D. and Constanza Salazar Ph.D. Header/footer gallery: creating and sustaining an online only art gallery. Internet Histories, 8(1–2):86–99, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2249363.

Silva:2023:DPM

[210] Erik Da Silva. Digital Pierrot Museum from Pristina to the Moon: an interview with Willred Dallto. Internet Histories, 8(1–2):100–113, 2023. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2023.2290894.

Cao:2023:ROS

[211] Tiancheng Leo Cao. Rethinking openness: a social constructivist approach to the promises of the new museology. Internet Histories, 8 (1–2):114–135, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2244341.

Arthur:2023:HOA

[212] Paul Longley Arthur, Lydia Hearn, Isabel Smith, and Nikos Koutras. How “open” are Australian museums? A review through the lens of copyright governance. Internet Histories, 8 (1–2):136–152, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2268375.

Arias:2023:FDI

[213] Maria Paula Arias. Framing digital identities through social media in museums. Internet Histories, 8 (1–2):169–187, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2023.2290939.

Wagner:2023:BRT

[214] Travis L. Wagner. Book review: The Two Revolutions: a History of the Transgender Internet, by Avery Dame-Griff, New York, New York University Press, 2023. 272 pp., ISBN 978-1-4798-1831-0, $30 (paperback), $89 (hardcover). Internet Histories, 8 (3):263–266, 2023. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2023.2299542.

deWild:2024:EMW

[215] Karin de Wild and Nadezhda Povroznik.

Editorial: Museums on the Web. Internet Histories, 8(1–2):1–8, 2024. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2024.2346870.

Tinker:2024:AAC

[216] Amanda J. Tinker. Accessing the artwork in COVID-19: loss, recovery and reimagination. Internet Histories, 8(1–2):46–61, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2306772.

Povroznik:2024:MDI

[217] Nadezhda Povroznik. Museums’ digital identity: key components. Internet Histories, 8(1–2):153–168, 2024. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2024.2333094.

Skov:2024:DCA

[218] Mette Skov and Tanja Svarre. A diachronic cluster analysis of Danish museum websites. Internet Histories, 8 (1–2):188–203, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2313292.

Mackinnon:2024:MWP

[219] Katie Mackinnon. The modem world: a pre-history of social media. Internet Histories, 8(1–2):204–207, 2024. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2024.2324406.

Whelan:2024:CFM

[220] Alice Whelan. Concealing for freedom: The making of encryption, secure messaging and digital liberties. Internet Histories, 8(1–2):208–212, 2024. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2024.2302292.

Passmann:2024:FUR

[221] Johannes Paßmann and Lisa Gerzen. Follow the updates! Reconstructing past practices with web archive data. Internet Histories, 8(3):213–228, 2024. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2024.2310405.

Angelou:2024:GLM

[222] Ioannis Angelou and Andreas Veglis. Greek legacy media organizations in the digital age: a historical perspective of web tool adoption (1990s–2023). Internet Histories, 8(3):229–245, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2024.2303697.

Tofalvy:2024:OPW

[223] Tamas Tofalvy. “The offline part was the essence of it” – the social history of iWiW, the Hungarian online social network. Internet Histories, 8(3):246–262, 2024. CODEN ????

ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2370748.

Smith:2024:BRH

[224] Alexander O. Smith. Book review: A History of Fake Things on the Internet, by Walter J. Scheirer, Stanford University Press, 2023, 264 pp, hardcover ISBN: 978-1-5036-3288-2, ebook ISBN: 978-1-5036-3704-7. Internet Histories, 8(3):267–270, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2346376.

Ma:2024:BRC

[225] Lai Ma. Book review: The Chinese Computer: a Global History of the Information Age, by Thomas S. Mullaney, The MIT Press, 2024, 376 pp, ISBN: 978-0-262-04751-7. Internet Histories, 8(3):271–273, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2352960.

Brugger:2024:TIH

[226] Niels Brügger, Anat Ben-David, Ian Milligan, and Valérie Schafer. The third internet histories Early Career Researcher Award. Internet Histories, 8(4):275–276, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2421583.

Brewster:2024:FRN

[227] Kat Brewster. “The flames are 50/ 50 right now”: content moderation practices at the onset of the HIV/ AIDS epidemic in the United States (1982–1990). Internet Histories, 8 (4):277–293, 2024. CODEN ????

ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2024.2416296.

Riemens:2024:FEW

[228] Rianne Riemens. Fixing the earth: whole-systems thinking in Silicon Valley’s environmental ideology. Internet Histories, 8(4):294–311, 2024. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2024.2416295.

Galip:2024:MEC

[229] Idil Galip. Methodological and epistemological challenges in meme research and meme studies. Internet Histories, 8(4):312–330, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2024.2359846.

West:2024:BRD

[230] Tyler West. Book review: The Digital Revolution: a Short History of an Ideology, By Gabriele Balbi, Oxford University Press, 2024, 176 pp., ISBN 978-0-19-887597-0. Internet Histories, 8(4):331–334, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2024.2360321.

Davoudi:2024:BRD

[231] Nilou Davoudi. Book review: The Digital Departed: How We Face Death, Commemorate Life, and Chase Virtual Immortality, By Timothy Recuber, NYU Press, 2023, 288 pp., ISBN: 978-1-4798-1494-7. Internet Histories, 8(4):335–339, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2024.2360320.

Pettis:2024:BRW

[232] Ben T. Pettis. Book review: WhatsApp: From a one-to-one Messaging App to a Global Communication Platform, By Amelia Johns, Ariadna Matamoros-Fernández, and Emma Baulch. 2024. Polity Press, 220pp. ISBN 978-1-5095-5052-4. Internet Histories, 8(4):340–343, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2380637.

Fortunati:2025:IRW

[233] Leopoldina Fortunati, Autumn Edwards, and Janet Abbate. Introduction: (re)writing gender in Internet histories. Internet Histories, 9(1–2):1–10, 2025. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2025.2458990.

Vincent:2024:SDG

[234] Jane Vincent. “She does a good job for a woman” — the challenges for women enablers in the era of Internet histories. Internet Histories, 9(1–2):11–30, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2424610.

Jankowski:2024:BWW

[235] Steve Jankowski. Becoming Wikipedian women: a sociotechnical history of the Gender Gap Task Force (2013– 2023). Internet Histories, 9(1–2):31–52, 2024. CODEN ???? ISSN 2470-1483.

URL http://www.tandfonline.com/
doi/full/10.1080/24701475.2024.
2425150.

Mari:2024:GIU

[236] Will Mari and Fanny Ramirez. Gendering of the “ideal user” in 1990sera internet magazines. Internet Histories, 9(1–2):53–68, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2024.2386760.

Kustritz:2024:PDT

[237] Anne Kustritz. The politics of the digital transition: lessons from slash fan fiction communities at the turn of the millennium. Internet Histories, 9(1–2):69–85, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2024.2442884.

Edwards:2024:DPM

[238] Emily Lynell Edwards. Digital pioneers: Mormon mommy bloggers and building the “Bloggernacle”. Internet Histories, 9(1–2):86–106, 2024. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2024.2438464.

Carradore:2024:SSS

[239] Roberto Carradore and Tiziana Pirola. Streaming out of the spiral of silence. Women’s visibility in gaming community and on twitch. Internet Histories, 9(1–2):107–127, 2024. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2024.2431800.

Brewster:2024:NBQ

[240] Kat Brewster. Network breakdown: the queer anarchist politics at the heart of the ’net from FidoNet to HOMOCORE. Internet Histories, 9 (1–2):128–143, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2424609.

Avenant:2024:SGS

[241] Shell Avenant. (Un)scaffolding gender: social media and the evolution of nonbinary identity. Internet Histories, 9(1–2):144–162, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2431797.

Melis:2024:HYM

[242] Beatrice Melis, Marta Fioravanti, Chiara Paolini, and Daniele Metilli. How have you modelled my gender? Reconstructing the history of gender representation in Wikidata. Internet Histories, 9(1–2):163–179, 2024. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2024.2431798.

Verhoef:2024:DLI

[243] Jesper Verhoef. Doing LGBTQ Internet histories justice: a queer web archive manifesto. Internet Histories, 9(1–2):180–197, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2357897.

MacDonald:2024:DFM

[244] Shana MacDonald, Brianna Wiens, and Nick Ruest. Dwelling with feminist media archives in the age of big data. Internet Histories, 9 (1–2):198–215, 2024. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2024.2431799.

Abbate:2024:RGH

[245] Janet Abbate, Cassius Adair, Avery Dame-Griff, Autumn Edwards, Leopoldina Fortunati, Deena Larsen, Laine Nooney, and Valérie Schafer. Researching gender in the history of the Internet and the Web. A roundtable at the SHOT 2023 conference. Internet Histories, 9(1–2):216–228, 2024. CODEN ???? ISSN 2470-1483. URL http:/

                                        /www.tandfonline.com/doi/full/10.
                                        1080/24701475.2024.2335023.

Coleman:2025:EWH

[246] Miles C. Coleman and Will Mari. An early Web history of vaccine skeptical digital rhetorics. Internet Histories, 9(3):229–251, 2025. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2025.2483568.

Kjellstrom:2025:PPS

[247] Zakayo Kjellström. Patterns of piracy: Sci-Hub and Sweden 2011–2018. Internet Histories, 9(3):252–273, 2025. CO-DEN ???? ISSN 2470-1483. URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/24701475.2025.2482459.

Stuckey:2025:BDA

[248] Helen Stuckey and Stephanie Harkin. Brilliant digital: a 1990s Australian videogames studio that brought Xena, KISS, and Popeye together in the Multipath Movies experimental streaming service. Internet Histories, 9 (3):274–291, 2025. CODEN ???? ISSN 2470-1483. URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        24701475.2025.2450177.

Owens:2025:PPC

[249] Trevor Owens, Benjamin Charles Germain Lee, and Jonah Estess. Powell.pps: close & distant reading of primary sources in web archives. Internet Histories, 9(3):292–311, 2025. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2025.2516306.

Hegarty:2025:ADD

[250] Kieran Hegarty. Averting the Digital Dark Age: How archivists, librarians, and technologists built the Web a memory. Internet Histories, 9(3):312–315, 2025. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/10.1080/
24701475.2025.2465221.

Gorham:2025:TDW

[251] Ashley E. Gorham. Tor: From the Dark Web to the future of privacy. Internet Histories, 9(3):316–318, 2025. CODEN ???? ISSN 2470-1483. URL http:/

/www.tandfonline.com/doi/full/10.
1080/24701475.2025.2465006.